# Can Agents Fix Agent Issues?

### v1

@@ 5.1 Experimental Setup
SE Agent LLM Plausibly resolvedCorrectly resolvedLocalization % Avg. File-level Function-level \$Cost Agentless GPT-4o 18.67 6.00 22.97 14.40 0.65 Claude-3.5-S 12.67 8.67 21.86 11.54 0.33 AutoCodeRover GPT-4o 12.67 4.67 17.05 11.30 0.23 Claude-3.5-S 17.33 12.67 25.61 18.67 0.05 SWE-agent GPT-4o 5.33 3.33 12.65 11.77 1.15 Claude-3.5-S 6.67 6.67 15.58 11.26 0.57
@@ 5.2 Quantitative Results
Category ResolvedSub-category ResolvedTool-related issues 3/9 (33.33%) Tool dependency issues 2/3 (66.67%) Tool implementation issues 1/3 (33.33%) Memory-related issues 1/8 (12.50%) Memory content errors 1/5 (20.00%) LLM operation issues 2/11 (18.18%) Model access misconfiguration 1/2 (50.00%) Prompt-related issues 1/2 (50.00%) Workflow issues 1/6 (16.67%) - - Utility issues 8/14 (57.14%) Utility implementation issues 2/6 (33.33%) Utility dependency issues 2/2 (100.00%) Utility configuration issues 4/6 (66.67%)
@@ Appendix D Experiment Statistical Significance
Table  6 presents the mean results of SE agents on AgentIssue-Bench, along with their corresponding two-sigma (±$\pm{}2\sigma$) errors. To calculate the two-sigma errors, we conducted the experiment on AgentIssue-Bench three times, computed the standard deviation for the results and multiplied it by 2, as shown in Equation 1:
@@
Table 6: Mean results with 2-sigma(±$\pm{}2\sigma$) errors of SE agents on AgentIssue-Bench
SE Agent LLM Plausibly resolvedCorrectly resolvedLocalization % File-level Function-level Agentless GPT-4o 18.67 (±$\pm{}2.31$) 6.00 (±$\pm{}4.00$) 22.97 (±$\pm{}4.74$) 14.40 (±$\pm{}3.23$) Claude-3.5-S 12.67 (±$\pm{}2.31$) 8.67 (±$\pm{}2.31$) 21.86 (±$\pm{}3.81$) 11.54 (±$\pm{}1.36$) AutoCodeRover GPT-4o 12.67 (±$\pm{}4.62$) 4.67 (±$\pm{}4.62$) 17.05 (±$\pm{}1.79$) 11.30 (±$\pm{}1.03$) Claude-3.5-S 17.33 (±$\pm{}4.62$) 12.67 (±$\pm{}2.31$) 25.61 (±$\pm{}3.18$) 18.67 (±$\pm{}6.67$) SWE-agent GPT-4o 5.33 (±$\pm{}4.62$) 3.33 (±$\pm{}4.62$) 12.65 (±$\pm{}5.06$) 11.77 (±$\pm{}6.94$) Claude-3.5-S 6.67 (±$\pm{}2.31$) 6.67 (±$\pm{}2.31$) 15.58 (±$\pm{}7.00$) 11.26 (±$\pm{}1.59$)

### v2

@@ 5.1 Experimental Setup
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % Avg. File-level Function-level \$Cost Agentless GPT-4o 18.67 6.00 22.97 14.40 0.65 Claude-3.5-S 12.67 8.67 21.86 11.54 0.33 AutoCodeRover GPT-4o 12.67 4.67 17.05 11.30 0.23 Claude-3.5-S 17.33 12.67 25.61 18.67 0.05 SWE-agent GPT-4o 5.33 3.33 12.65 11.77 1.15 Claude-3.5-S 6.67 6.67 15.58 11.26 0.57
@@ 5.2 Quantitative Results
Category Resolved% Sub-category Resolved% Tool-related issues 3/9 (33.33%) Tool dependency issues 2/3 (66.67%) Tool implementation issues 1/3 (33.33%) Memory-related issues 1/8 (12.50%) Memory content errors 1/5 (20.00%) LLM operation issues 2/11 (18.18%) Model access misconfiguration 1/2 (50.00%) Prompt-related issues 1/2 (50.00%) Workflow issues 1/6 (16.67%) - - Utility issues 8/14 (57.14%) Utility implementation issues 2/6 (33.33%) Utility dependency issues 2/2 (100.00%) Utility configuration issues 4/6 (66.67%)
@@ Appendix D Experiment Statistical Significance
Table  6 presents the mean results of SE agents on AgentIssue-Bench, along with their corresponding two-sigma ($\pm{}2\sigma$) errors. To calculate the two-sigma errors, we conducted the experiment on AgentIssue-Bench three times, computed the standard deviation for the results and multiplied it by 2, as shown in Equation 1:
@@
Table 6: Mean results with 2-sigma($\pm{}2\sigma$) errors of SE agents on AgentIssue-Bench
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 18.67 ($\pm{}2.31$) 6.00 ($\pm{}4.00$) 22.97 ($\pm{}4.74$) 14.40 ($\pm{}3.23$) Claude-3.5-S 12.67 ($\pm{}2.31$) 8.67 ($\pm{}2.31$) 21.86 ($\pm{}3.81$) 11.54 ($\pm{}1.36$) AutoCodeRover GPT-4o 12.67 ($\pm{}4.62$) 4.67 ($\pm{}4.62$) 17.05 ($\pm{}1.79$) 11.30 ($\pm{}1.03$) Claude-3.5-S 17.33 ($\pm{}4.62$) 12.67 ($\pm{}2.31$) 25.61 ($\pm{}3.18$) 18.67 ($\pm{}6.67$) SWE-agent GPT-4o 5.33 ($\pm{}4.62$) 3.33 ($\pm{}4.62$) 12.65 ($\pm{}5.06$) 11.77 ($\pm{}6.94$) Claude-3.5-S 6.67 ($\pm{}2.31$) 6.67 ($\pm{}2.31$) 15.58 ($\pm{}7.00$) 11.26 ($\pm{}1.59$)